# Mixed-Integer Programming STALC Stratified Topological Autonomy for Long-Range Coordination

### v1

@@ IX-A2 Bounding Overwatch Example
Figure 10: Example solution demonstrating bounding overwatch behavior as all robots move from node 1 to 11. Robot team routes are shown with each time step labeled. The overwatch opportunities, indicated by the pink arrows, are met at time steps 2, 3, 6, and 7. The two teams alternate providing overwatch as they move through the environment.
As a demonstration of a larger graph that particularly lends itself to the bounding overwatch paradigm, Fig. 10 shows a solution to our MIP problem for the graph shown in blue. We set the start point for all robots to be node 1 and the destination for all robots is node 11. The solution yields two robot teams alternating traversing and providing overwatch. Each team’s route is visualized along the edges with time steps noted to enable correlating the teams’ positions and the bounding overwatch behavior.
@@ IX-A3 Meadow Map 1
In Figures 11 and 12, we demonstrate applying this approach to complex real world scenarios. We consider traversing between areas of forested cover in a high-risk environment, using a simulated meadow environment from [41] as a representative example. We show results with all robots reaching a designated goal and a subset of robots reaching the goal.
In Fig. 11, we consider 10 robots starting at node 1 with a goal of all robots reaching node 2 within 10 time steps. In our generated solution, three distinct paths emerge. Through time step 3, six robots maneuver together (orange and red teams) alternating providing overwatch with another team (purple) of four robots. At time step 4, two robots (red) remain at node 4 to provide overwatch for the purple team, while orange maneuvers toward the next overwatch position. This allows the purple and red teams to combine and proceed alternating providing overwatch with the orange team through the rest of the graph toward the goal node.
Figure 12: Reconnaissance scenario spanning a meadow environment. The nodes are placed in forested regions of cover. Ten robots start at node 1 with the goal of at least one robot reaching node 2, across the meadow. The edges of the graph have cost for transitioning between nodes that is a function of their distance, detectability, and vulnerability. This cost can be reduced by moving in teams and by providing overwatch, where robots at a particular node oversee the movement of robots along a corresponding edge to help mitigate some of the risk of traversing. Overwatch opportunities are indicated with an arrow from the overwatch node pointing toward the edge that can be monitored. The robot team routes through the graph, solved for with our proposed method, represent a solution to this problem.
@@ IX-A4 Meadow Map 2
Fig. 12 is an example with further subdivision of the robots in the solution. Again, we consider 10 robots starting at node 1. In this scenario, the goal is for at least one robot to reach node 2 within 12 time steps. In this case, the solution has four teams emerge. One team of four (orange) that traverses through the map towards the goal, providing overwatch as needed, and three smaller supporting teams (red, purple, brown) providing overwatch for orange and each other. Since all robots do not need to reach the goal node, this example highlights the trade-off between the cost of traversing and the benefit of overwatch. Teams will not continue towards the goal if the overall cost of their movement is not less than the benefit of the overwatch they would provide to a primary team moving towards the goal. This concept is consistent with expectations in an operational setting: the risk outweighs the reward. Ultimately, this is the objective of our method, to be able to determine strategic tactical maneuvers in complex environments.
@@ IX-A5 Ablation Study
Figure 13: Ablation study illustrating the impact of each of our main constructs: overwatch opportunities, edges with high vulnerability, and incentives for moving as a team. Robot team paths are shown in each subplot with components of our formulation incrementally added.
We performed an ablation study to assess the impact of the constructs we present in Sec. III, which we label overwatch, vulnerability, and teaming, in our algorithm. Fig. 13 shows the solutions to our illustrative example when removing these components and incrementally adding them back in. The scenario is setup as in Fig. 9, with all robots starting at node 1 and an overall goal of at least one robot reaching node 5.
When we remove the overwatch, vulnerability, and teaming constructs, in Fig. 13(a), the problem essentially becomes a shortest path problem. The weights in the graph are fixed and one robot takes the path with the least overall cost due to the edge weights and time. It would be an equivalent cost solution for any number of robots to take this path.
In tactical maneuvers, overwatch allows minimizing detection and maximizing safety. When we add our formulation of overwatch opportunities in Fig. 13(b), the edge weights now vary based on the overwatch opportunities being utilized by the team and the optimal solution includes robots moving to the two overwatch positions while one robot traverses to the goal, overall making the operation safer.
As an incentive to travel in a formation on edges that are particularly dangerous, or vulnerable, we add our vulnerability construct back in Fig. 13(c), making edges (1,3), (3,5), and (4,5) more costly to traverse alone. We see in the solution that the desired minimum of 4 robots traverse edge (4,5). The robot providing overwatch at node 3 moves with other robots on edge (1,2) and then alone on edge (2,3). This is ultimately the same cost as if four robots were to traverse the vulnerable edge (1,3). Both solutions are optimal. As in the previous cases, the extra robots stay at the start node.
In a tactical scenario, more robots moving together would provide more data, potentially reducing the overall threat level. Additionally, the redundancy of having more robots yields a more robust solution. When we add our teaming construct back into the problem, as in Fig. 13(d), we see all robots moving for the first time, no longer leaving robots at the start node, since moving as a team provides cost reductions on each edge and greater cost reductions when providing overwatch. Ultimately, the overwatch, vulnerability, and teaming constructs yield practical plans for tactical maneuvers that minimize detection and maximize robustness and safe navigation.
@@ IX-B Graph Planning Computation Time
The common limitation of MIP is computation time. The high level planning we propose would need to be computed rapidly (on the order of magnitude of minutes) for robots to be able to act on the plan and re-plan as new data is gathered. This allows a new plan to be generated while the robots traverse one edge of about $\sim 100-250m$. Table IV shows the parameters that affect the problem size for the graphs discussed in this paper (number of locations, number of overwatch opportunities, and time horizon), the total variables in our optimization problem due to those parameters, and the resulting solve time for our MIP problem in each case. The graphs are labeled Illustrative, Bounding, Map 1, and Map 2, which correspond to Figures 9 and 13, Fig. 10, Fig. 11, and Fig. 12, respectively.
(a) Meadow environment [41]
(b) Visibility map
(c) Pruned A paths
(d) Node 1 overwatch visibility
(e) Dynamic topological graph
(f) Robot team path solution
@@
Figure 15: Environment segmentation for two different observer locations in an forested off-road environment. Resulting cover regions and A paths between the regions are depicted for each.
We used the Gurobi optimizer [39] on an Intel® Core™ i7-10875H CPU @ 2.30GHz × 16 and averaged the solve times across 100 trials. In all cases, the number of robots was 10. Varying the number of robots does not impact the total number of variables in our problem. We select suitable time horizons for each problem based on the size of the graph and location of the goal to ensure that the problem is feasible.
These results show computation times in seconds for realistic operational scenarios. Additionally, the problem stays computationally tractable to solve complicated routes on large graphs for multi-robot teams.
IX-C STALC Simulated Environment Segmentation and Graph Planning
@@
Figure 16: Schematic forested graphs showing overwatch opportunities and team routes for the denoted start and goal nodes. Both robots start at node 0 and the goal is for one robot to reach each goal node (nodes 6 and 7). Two different weights are used for Forested Graph 1 to show the difference in the resulting paths.
@@ IX-D STALC Hardware Experiments
We evaluated our hierarchical planner STALC on hardware to experimentally validate our graph generation and visibility calculations for real-world environments. Additionally, we wanted to evaluate operating with a decoupled high-level and low-level planning paradigm to ensure that the high-level plan is executable on a real-world system.
We demonstrated the ability to achieve minimum visibility planning with a team of autonomous ground vehicles with our hierarchical approach: by coupling our high-level graph-based planner with our mid-level A planner and low-level stochastic model predictive control approach. To do this, we constructed both a visibility map and a traversability map from a priori data the environment.
We evaluated our approach in two highly constrained environments: forested off-road terrain and an urban environment. In both cases, the vehicles needed to avoid both static and dynamic obstacles (team members), operate in formation, and provide overwatch to their team members. We considered two different observer locations in both environments to demonstrate how the environment segmentation and ultimate route through the environment will vary based on the scenario. Additionally, we show results with different emphases on overwatch versus moving in formation to demonstrate the adaptability of our graph planner to different scenario objectives.
@@
Figure 18: Overwatch view of robot Wanda while robot Willow traverses an edge. Observer view on the top left shows the traversing robot being detectable from the observer perspective.
Figure 19: Depiction of the route the robots traversed Forested Graph 1 with the regions the observer could see the robots highlighted.
@@ IX-D1 Forested Off-road Environment
We consider a trapezoidal region of a forested off-road environment with varying types of brush, trees, and terrain. We segment the environment around two different observer locations, as depicted in Fig. 15. The change in the observer location results in significant variations in the cover regions and A paths, ultimately resulting in different topologies for our dynamic graphs.
@@
(a) Urban Graph 1 with overwatch scaled by 1
(b) Urban Graph 1 with overwatch scaled by 1.3
Figure 21: Planned routes on Urban Graph 1 showing the difference in routes when the overwatch scale increases. In this case, this change resulted in different overwatch opportunities and a significant difference in the resulting team paths through the graph.
(a) Urban Graph 2 with overwatch scaled by 1
(b) Urban Graph 2 with overwatch scaled by 0.4
Figure 22: Planned routes on Urban Graph 2 showing the difference in routes when the overwatch scale decreases. In this case, this change resulted in different overwatch opportunities and a subtle change to the resulting team paths through the graph.
@@
In Fig. 16, we illustrate a schematic representation of our topological graph with the overwatch opportunities indicated. We depict the optimal team paths through the graph along the edges, as solved for using the algorithm in Table III. Between Fig. 16(a) and Fig. 16(b), we varied the value of overwatch by allowing more overwatch opportunities to be considered in Fig. 16(b). We do this by setting the allowable ratio of overwatch reduction to edge cost (as described in Sec. V-C) from 50% to 10%. When overwatch opportunities are available, they offer a significant reduction to the edge costs and thus are desirable to be used. This results in the robots moving in formation on edge (3,6) when overwatch is less valuable versus performing bounding overwatch in the other case. Ultimately, in an operational scenario these types of factors and thresholds would be set based on the priorities of the scenario and the perceived risk associated with performing overwatch compared to formation maneuvers (i.e., if a small amount of overwatch would still be valuable to consider versus ignored in favor of moving as a team). These two different solutions on the same graph demonstrate the adaptivity of our algorithm to varying objectives.
Fig. 16(b) and Fig. 16(c) utilize the same graph planner parameters on the graphs generated from the different observer locations in Fig. 15. The observer location significantly impacts the formulation of the dynamic topological graph and overwatch opportunities, ultimately resulting in significantly different routes through the environment.
Fig. 17 depicts the robots moving in formation on edge (3,6), following the paths in Fig. 16(a). The vehicles maintain a set following distance along the A paths while avoiding environmental obstacles.
Fig. 18 shows the robot Wanda performing overwatch at node 3 while the robot Willow traverses along edge (0,3). Each robot is equipped with a 360 degree camera. Additionally, the top left corner shows the observer view while Wanda is visible from the observer’s perspective.
The route traversed by the robots in Forested Graph 1, in Fig. 16(a) and Fig. 16(b), is outlined in Fig. 19. Additionally, Fig. 19 highlights the regions of this route where the robots were observed by a camera located at the observer position. This demonstrates that our cover regions were a conservative estimate of areas that would not be visible from the observer location.
The included video shows the robot team traversing the off-road environment using STALC following the routes in Fig. 16(a).
@@ IX-D2 Protection Metric
To evaluate the experimental performance of the multi-robot team, we propose the following metric to characterize the level of protection of the traversing team.
$\mathcal{M}_{protect}=\frac{1}{n_{A}}\sum_{a=1}^{n_{A}}\frac{1}{D_{a}}\sum_{e\in E}(d_{o,ae}+d_{f,ae}+d_{c,ae})$ (46)
In this metric, we consider for each robot $a$ and edge $e$, the distance a robot was monitored through overwatch $d_{o,ae}$, the distance traversed in formation $d_{f,ae}$, and the distance traversed in cover $d_{c,ae}$, as these are the three different forms of protection the robots can utilize for safe traversal of a region. The total distance traveled by a particular robot is $D_{a}$. We evaluate the protection metric to compare the level of protection between different scenarios. When a vehicle utilizes one form of protection (traversing in cover, formation, or being overwatched) throughout the entire path then this metric is $\mathcal{M}_{protect}=1$. Additional protection will increase this score to a maximum of $3$.
We consider three scenarios on Forested Graph 1 to compare the metric: two robots moving in formation on edge (3,6) following the plan in Fig. 16(a), two robots moving in bounding overwatch from Fig. 16(b), and a single robot traversing the route shown in Fig. 19. We present the value of the protection metric for each of these cases in Table V. As expected, the single robot scenario has the lowest score since it is not able to be further protected by a teammate with bounding overwatch or moving in formation. In this scenario, the robots moving in formation on edge (3,6) received a higher protection metric than bounding overwatch since they can maintain visual contact with each other the entire path. In comparison, the duration of overwatch on that edge is shorter due to the brush and other terrain blocking the robot’s field of view. However, the protection metric could be weighted appropriately if one form of protection is determined to be superior to the other for a particular scenario. For each of these scenarios, we computed the protection metric using camera data. A vehicle is considered to be in cover when it is not visible from the observer’s location.

### v2

@@ IX-A2 Bounding Overwatch Example
Figure 9: Example bounding overwatch solution as all robots move from node 1 to $1$. Robot coalition routes are displayed with the position of each coalition at each time step denoted by square markers along the routes. Time steps are labeled sequentially from $t0$ to $t9$. The overwatch opportunities, indicated by the pink arrows, are met at time steps 2, 3, 6, and 7. The two coalitions alternate providing overwatch as they move through the environment.
As a demonstration on a graph that particularly lends itself to the “bounding overwatch” paradigm, where robots alternate providing overwatch, Fig. 9 shows a solution to our MILP problem for the graph shown in blue. The robots start at node 1 and need to reach node 11. The solution yields two robot teams alternating traversing and providing overwatch. In this figure and subsequent figures, we depict our solution to the problem through routes for each coalition of robots. Along these routes, we denote each time step with square markers and text indicating the sequential step in the plan that the coalition traverses the corresponding edge or stops at the corresponding node. These markers enable correlating the relative positions of the coalitions’ to identify when coalitions are moving together or overwatch is occurring (i.e., one coalition is at an overwatch node while the other one traverses the corresponding edge at the same time step).
@@ IX-A3 Meadow Map 1
In Figures 10 and 11, we demonstrate traversing between areas of cover in a high-risk environment, using a simulated meadow environment from [65].
In Fig. 10, we consider 10 robots starting at node 1 with a goal of all robots reaching node 2 within 10 time steps. In our generated solution, three distinct paths emerge. Through time step 3, six robots maneuver together (orange and red teams) alternating providing overwatch with another team (purple) of four robots. At time step 4, two robots (red) remain at node 4 to provide overwatch for the purple team, while orange maneuvers toward the next overwatch position. This allows the purple and red teams to merge and proceed alternating providing overwatch with the orange team through the rest of the graph toward the goal node.
Figure 11: Reconnaissance scenario spanning a meadow environment. The nodes are placed in forested regions of cover. Ten robots start at node 1 with the goal of at least one robot reaching node 2, across the meadow. The robot team routes through the graph, solved for with our proposed method, show extensive utilization of overwatch opportunities.
@@ IX-A4 Meadow Map 2
Fig. 11 is an example with further subdivision of the robots in the solution. Again, we consider 10 robots starting at node 1. In this scenario, the goal is for at least one robot to reach node 2 within 12 time steps. In this example, four coalitions emerge: a primary coalition of four robots (orange) advancing toward the goal while providing overwatch as needed, and three smaller coalitions (red, purple, brown) supporting the primary coalition and each other. Since all robots do not need to reach the goal node, this example highlights the trade-off between the cost of traversing and the benefit of overwatch. The supporting coalitions eventually halt their movement, as the cost of traversing does not outweigh the benefits of overwatch they could provide. This concept is consistent with expectations in an operational setting: the risk outweighs the reward. Ultimately, our approach’s objective is to determine tactical maneuvers in complex environments.
@@ IX-A5 Ablation Study
Figure 12: Ablation study illustrating the impact of key components in our algorithm: overwatch opportunities, edges with high vulnerability, and incentives for moving as a team. Robot team paths are shown in each subplot with components of our formulation incrementally added.
(a) 5 Node Graphs
(b) 15 Node Graphs
(c) 25 Node Graphs
(d) 50 Node Graphs
Figure 13: Comparison of average computation time across state-of-the-art algorithms for multi-robot planning with overwatch opportunities across various graph sizes, edge densities, and numbers of robots. Each algorithm is adapted to solve the same problem and our approach is labeled as “MILP.”
Figure 14: Optimality gap versus the number of robots for the learning-based PPO algorithm. All other approaches were solved to optimality.
We performed an ablation study to assess the impact of providing overwatch, formations on vulnerable edges, and moving as a team wherever possible. Fig. 12 shows the solutions to our illustrative example, from Fig. 8, when ablating these components.
When we remove the overwatch, vulnerability, and teaming components, in Fig. 12(a), the problem becomes a shortest path problem. The weights in the graph are fixed and one robot takes the path with the least overall edge and time cost.
In tactical maneuvers, overwatch allows minimizing detection risk. When we add our formulation of overwatch opportunities in Fig. 12(b), the edge weights now vary based on the overwatch opportunities being utilized by the team and the optimal solution includes robots moving to the two overwatch positions while one robot traverses to the goal.
As an incentive to travel in a formation on edges that are particularly dangerous, or vulnerable, we add our vulnerability component back in Fig. 12(c), making edges (1,3), (3,5), and (4,5) more costly to traverse alone. We see in the solution that the desired minimum of 4 robots traverse edge (4,5).
In a tactical scenario, more robots moving together would be advantageous (e.g., due to redundancy or additional sensing). When we add our teaming component back into the problem in Fig. 12(d), we see all robots moving for the first time, no longer leaving robots at the start node, since moving as a team provides cost reductions on each edge and greater cost reductions when providing overwatch. Ultimately, the overwatch, vulnerability, and teaming components yield practical plans for tactical maneuvers that minimize detection and risk.
(a) Meadow environment [65]
(b) Visibility map
(c) Pruned A paths
(d) Node 1 overwatch visibility
(e) Dynamic topological graph
(f) Robot team path solution
@@ IX-B Graph Planning Computation Time
IX-B1 Comparison to State-of-the-Art (SOTA)
We compare our MILP approach to the current SOTA baselines for multi-robot planning with support: (i) JSG, which enumerates a joint state graph for all combinations of robots at locations and then performs a shortest path search [19]; (ii) PPO, a reinforcement learning approach that trains on a single graph and targets the challenge of scaling to larger graphs and teams [33]; and (iii) GMIP, a generic MIP formulation, based on standard paradigms [25, 55], that enumerates binary variables for each robot’s location at each time step (or $n_{T}n_{L}n_{A}$ variables). We adapted each algorithm to solve the same underlying problem. For this purpose, we compare to a simplified version of our algorithm (excluding the complexity of vulnerability and teaming), which we label “MILP,” to show the benefits of our formulation. In particular, we aim to show the major impact of removing the dependence on the number of agents in our decision space (scaling by $n_{T}n_{L}$ variables).
We evaluate each algorithm for graph sizes from 5 to 50 nodes with edge density levels: “Sparse,” “Moderate,” and “Dense,” which correspond to 20%, 50%, and 80% of a fully connected graph’s edges. We set 40% of those edges to have up to 2 overwatch opportunities. We then solve for multi-robot plans with team sizes in the set $\{2-10,50,100,200\}$. Start and goal positions are selected based on the furthest apart nodes by edge cost and the MIP-based approaches use a time horizon twice the shortest-path length.
PPO is trained per graph, so we compare training and inference time to JSG’s construction and shortest path search time, as in [33]. For GMIP and our MILP approach, we present the computation time to an optimal solution using the Gurobi optimizer [60]. All algorithms were run on a 13th Gen Intel® Core™ i9-13950HX × 32 CPU and PPO additionally used a NVIDIA GeForce RTX 4090 Laptop GPU. We averaged computation times across three seed values for each graph and consider failure when, for any of the seed values, the solve time exceeds 30 minutes, requires more than 16GB of RAM, or (in the case of PPO) does not successfully reach the goal. We terminate training for PPO when the rolling 50-episode average reward changes by less than 0.2 for 500 episodes, as in [33].
Fig. 13 compares the average computation time of each algorithm for four graph sizes versus the number of robots. The computation time and number of robots are shown on a log-scale with the same Y-axis used in each plot to see relative magnitude. In these plots, less computation time is ideal. Rapid computation time (on the order of magnitude of minutes for edge lengths of about 100250m100-250m) is essential to enable re-planning with new information. Our MILP approach finds optimal solutions in under 5 seconds across all graph sizes, densities, and team sizes. This is several orders of magnitude faster than the baselines, which quickly become intractable for larger numbers of robots and graph sizes. Notably, in some cases, the solve time of our MILP decreases for larger number of agents since the problem becomes easier when more overwatch opportunities can be utilized.
JSG, GMIP, and MILP all return optimal solutions. In Fig. 14, we show the optimality gap for the PPO approach relative to our MILP formulation. In this plot, a larger ratio is better (i.e., 1.0 represents optimality). PPO’s optimality ratio decreases with team size and graph density, highlighting PPO’s decreasing solution quality in these regimes.
IX-B2 Example Graphs’ Computation Time
We additionally report the computation time for the simulated graphs discussed in this paper, where we consider a more complex problem with teaming and vulnerability considerations. Table IV shows the parameters that affect the problem size (i.e., $n_{L},n_{O},n_{T}$), the resulting total optimization variables, and the solve time for our graph planner averaged across 100 trials. In all cases, the number of robots was 10. Varying the number of robots does not impact the total number of variables in our problem. The graphs are labeled Illustrative, Bounding, Map 1, and Map 2, which correspond to Fig. 8, Fig. 9, Fig. 10, and Fig. 11, respectively. We select time horizons for each problem based on the size of the graph and location of the goal to ensure that the problem is feasible. These results show computation times in seconds for complex operational scenarios.
@@ IX-B2 Example Graphs’ Computation Time
Figure 16: Graph generation for two different observer locations in an forested off-road environment. Resulting cover regions and A paths between the regions are depicted for each.
IX-C STALC Simulated Graph Generation and Planning
@@
Figure 17: Schematic forested graphs showing overwatch opportunities and team routes. Both robots start at node 0 and the goal is for one robot to node 6 and one robot to reach node 7. The priority of overwatch is varied in Forested Graph 1 to show the difference in the resulting paths.
@@ IX-D STALC Hardware Experiments
We evaluated our hierarchical planner STALC in hardware experiments to validate our graph generation and visibility calculations in real-world environments. Additionally, we evaluated the feasibility solving (1) using a high-level and low-level planning paradigm. We used Clearpath Warthogs in these experiments as a surrogate platform, though our hierarchical planner is broadly applicable to other systems.
Our hardware experiments demonstrate minimum-visibility planning with a team of autonomous ground vehicles with our hierarchical approach: coupling our high-level graph-based planner with our mid-level A planner and low-level stochastic NMPC. We evaluated our approach in two highly constrained environments: forested off-road terrain and an urban environment. In both cases, the vehicles needed to avoid both static and dynamic obstacles (coalition members), operate in formation, and provide overwatch to their team members. We use a priori data of the environment to construct our visibility and traversability maps. We considered two different observer locations in each environment to demonstrate how the environment segmentation and ultimate route through the environment will vary based on the scenario. Additionally, we show results with different emphases on overwatch versus moving in formation to demonstrate the adaptability of our graph planner to different scenario objectives.
@@
Figure 19: Overwatch view of robot Wanda while robot Willow traverses an edge. Observer view in the top left corner shows the traversing robot being detectable from the observer perspective.
Figure 20: Depiction of the route traversed by the robots in Forested Graph 1, with path segments visible to the observer highlighted.
@@ IX-D1 Forested Off-road Environment
We consider a trapezoidal region of a forested off-road environment with varying types of brush, trees, and terrain. We segment the environment around two different observer locations, as depicted in Fig. 16. The change in the observer location results in significant variations in the cover regions and A paths, ultimately resulting in different topologies for our DTG.
@@
(a) Urban Graph 1: 1.0 overwatch scale
(b) Urban Graph 1: 1.3 overwatch scale
Figure 22: Planned routes on Urban Graph 1 showing the difference in routes when the overwatch scale increases. In this graph, the overwatch scale change resulted in different overwatch opportunities and a significant difference in the resulting team routes through the graph.
(a) Urban Graph 2: 1.0 overwatch scale
(b) Urban Graph 2: 0.4 overwatch scale
Figure 23: Planned routes on Urban Graph 2 showing the difference in routes when the overwatch scale decreases. In this graph, the overwatch scale change resulted in different overwatch opportunities and a subtle change to the resulting team paths through the graph.
@@
In Fig. 17, we depict our schematic representation of the corresponding topological graphs with the overwatch opportunities indicated and optimal team routes, as solved for using our graph planning approach. Between Fig. 17(a) and Fig. 17(b), we varied the priority of overwatch, which results in more overwatch opportunities in Fig. 17(b). We do this by setting the allowable ratio of overwatch reduction to edge cost (as described in Sec. VI-C) from 50% to 10%. When overwatch opportunities are available, they offer a significant reduction to the edge costs and thus are desirable to be used. This change results in the robots moving in formation on edge (3,6) when overwatch is less valuable versus performing bounding overwatch across that edge when overwatch is more valuable. Ultimately, in an operational scenario these types of factors and thresholds would be set based on the priorities of the scenario and the perceived risk associated with performing overwatch (i.e., the relative value of a small amount of overwatch compared to moving as a team). These two different solutions on the same graph demonstrate the adaptivity of our algorithm to varying objectives.
Fig. 17(b) and Fig. 17(c) utilize the same overwatch parameters on the graphs generated from the different observer locations in Fig. 16. The observer location significantly impacts the resulting DTG, and thus the routes through the environment.
Fig. 18 depicts the robots moving in formation on edge (3,6), following the paths in Fig. 17(a). The vehicles maintain a set following distance along the A paths while avoiding environmental obstacles. Fig. 19 shows the robot Wanda performing overwatch at node 3 while the robot Willow traverses along edge (0,3). Each robot is equipped with a 360 degree camera. Additionally, the top left corner shows the observer view while Wanda is visible from the observer’s perspective.
The route traversed by the robots in Forested Graph 1, in Figures 17(a) and 17(b), is outlined in Fig. 20. In Fig. 20, we highlight the regions of this route where the robots were observed by a camera located at the observer position. This demonstrates that our cover regions were a conservative estimate of areas that would not be visible from the observer location.
@@ IX-D2 Protection Metric
To evaluate the experimental performance of the multi-robot team, we propose the following metric to characterize the team’s level of protection:
$\mathcal{M}_{\textrm{protect}}=\frac{1}{n_{A}}\sum_{i=1}^{n_{A}}\frac{1}{D^{i}}\sum_{e\in E}(d^{i}_{o,e}+d^{i}_{f,e}+d^{i}_{c,e}).$ (39)
In this metric, for each robot $i$ and edge $e$, we consider the distances the robots traverse using three forms of protection: (i) the distance a robot was monitored through overwatch $d^{i}_{o,e}$; (ii) the distance traversed in formation $d^{i}_{f,e}$; and (iii) the distance traversed in cover $d^{i}_{c,e}$. The total distance traveled by a particular robot is $D^{i}$. When a vehicle utilizes one form of protection (traversing in cover, formation, or being overwatched) throughout the entire path, this metric is $\mathcal{M}_{\textrm{protect}}=1$. Additional protection will increase this score to a maximum of $3$.
We consider three scenarios on Forested Graph 1 to compare the levels of protection: (1) two robots moving in formation on edge (3,6) following the plan in Fig. 17(a), (2) two robots moving in bounding overwatch from Fig. 17(b), and (3) a single robot traversing the route shown in Fig. 20. For each of these scenarios, we computed the protection metric using camera data. A vehicle is considered to be in cover when it is not visible from the observer’s location. We present the protection metric values in Table V. As expected, the single robot scenario has the lowest score since it is not able to be further protected by a teammate. In this graph, the robots moving in formation on edge (3,6) received a higher protection metric than the bounding overwatch scenario since the robots in formation maintain visual contact with each other throughout the path. In comparison, the duration of overwatch on edge (3,6) is shorter due terrain features blocking the robot’s field of view. If one form of protection were determined to be superior to the others for a particular scenario, the protection metric could be weighted accordingly.